In a right triangle, the legs are 3 and 3 sqrt3. Find the length of the hypotenuse.
Answer by venugopalramana(3286)

In a right triangle, the legs are 3 and 3 sqrt3. Find the length of the hypotenuse.
HYPOTENUSE^2=3^2+{3SQRT(3)}^2=9+9*3=9+27=36
HYPOTENUSE =6